OpenStudy (help_people):

Samuel bought a cement mixer for $54,205. The value of the cement mixer depreciated at a constant rate per year. The table below shows the value of the cement mixer after the first and second years: Year 1 2 Value (in dollars) 47,158.35 41,027.76 Which function best represents the value of the cement mixer after t years? f(t) = 47,158.35(0.87)t f(t) = 54,205(0.13)t f(t) = 47,158.35(0.13)t f(t) = 54,205(0.87)t

Nnesha (nnesha):

1)post a question 2) tag :-)

OpenStudy (help_people):



Nnesha (nnesha):

depreciated meaning what ? any idea how to start ?

Nnesha (nnesha):

key words depreciated constant rate *per* year

OpenStudy (help_people):

multiply something but no idk where to start

Nnesha (nnesha):

depreciate =decrease

Nnesha (nnesha):

so \[\huge\rm f(t)=P (1- \frac{ r }{ 100 })^t\] r=rate p =starting amount t= years substitute values :-) and then sovle

OpenStudy (help_people):

where are the vlues?

Nnesha (nnesha):

read the question :-)

OpenStudy (help_people):

f(t)=54,205(1-r/100)^1?

Nnesha (nnesha):

alright great can you solve for r :-) and yeah when t=1 f(t) = 47158.35 :-)

Nnesha (nnesha):

\[47158.35 =54,205(1-\frac{r}{100})^1\]

OpenStudy (help_people):

you would diode 54,205 on the other side right

Nnesha (nnesha):

yep divide**

OpenStudy (help_people):

47185/54205=(1-6/100)^1

OpenStudy (help_people):

6 is r srry

Nnesha (nnesha):

yep right :-) solve for r

OpenStudy (help_people):

now I'm confused how to do that can u show me please?

Nnesha (nnesha):

47185/54205= ??

Nnesha (nnesha):

anything to the 1 power equal to same thing so (1-r/100)^1 = 1-r/100

OpenStudy (help_people):

-1?

Nnesha (nnesha):

nope 47185/54205= ?? divide this first

OpenStudy (help_people):

ok

OpenStudy (help_people):

0.87